OpenStudy (anonymous):

can someone give me a formula for this please? A college student earned $8400 during summer vacation working as a waiter in a popular restaurant. Part was invested at 8% simple interest and the remainder at 9% simple interest. At the end of one year, the student had earned $707 interest. How much was invested at 9%

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

707=0.08(8400 +0.09(8400)

OpenStudy (anonymous):

not possible

OpenStudy (amistre64):

the formula(s) would be: a + b = 8400; a = 8400-b 707 = a(.08) + b(.09) ; substitue the 'value' of a into this equation and solve for 'b'

OpenStudy (anonymous):

oops yes it is sorry. call the amount invested at 9% x. then \.09x+.08(8400-x)=707\]

OpenStudy (anonymous):

\[.09x+.08(8400-x)=707\] \[.01x+672=707\] \[.01x=35\] \[x=\frac{35}{.01}=3500\]